# Surg-3M: A Dataset and Foundation Model for Perception in Surgical Settings

### v2

@@ Abstract
Advancements in computer-assisted surgical procedures heavily rely on accurate visual data interpretation from camera systems used during surgeries. Traditional open-access datasets focusing on surgical procedures are often limited by their small size, typically consisting of fewer than 100 videos with less than 100K images. To address these constraints, a new dataset called Surg-3M has been compiled using a novel aggregation pipeline that collects high-resolution videos from online sources. Featuring an extensive collection of over 4K surgical videos totaling 938 hours of high-quality footage across multiple procedure types, Surg-3M offers a comprehensive resource surpassing existing alternatives in size and scope, including two novel tasks. To demonstrate the effectiveness of this dataset, we present SurgFM, a self-supervised foundation model pretrained on Surg-3M that achieves impressive results in downstream tasks such as surgical phase recognition, action recognition, and tool presence detection. Combining key components from ConvNeXt, DINO, and an innovative augmented distillation method, SurgFM exhibits exceptional performance compared to specialist architectures across various benchmarks. Our experimental results show that SurgFM outperforms state-of-the-art models in multiple downstream tasks, including significant gains in surgical phase recognition (+8.9pp, +4.7pp, and +3.9pp of Jaccard in AutoLaparo, M2CAI16, and Cholec80), action recognition (+3.1pp of mAP in CholecT50) and tool presence detection (+4.6pp of mAP in Cholec80). Moreover, even when using only half of the data, SurgFM outperforms state-of-the-art models in AutoLaparo and achieves state-of-the-art performance in Cholec80. Both Surg-3M and SurgFM have significant potential to accelerate progress towards developing autonomous robotic surgery systems. Our dataset, models, and code are publicly available at https://github.com/visurg-ai/surg-3m.
@@ 1 Introduction
Figure 1: Performance comparison between our foundation model, SurgFM, and the state-of-the-art (SotA) models. Our evaluation focuses on three surgical downstream tasks and six datasets. SurgFM results are shown in bold. Axis labels are written in regular font. For a detailed breakdown of the results, see Tables 3, 4, and 5.
Figure 2: Diversity and procedure prevalence in Surg-3M. Representative samples from various procedures, demonstrating the diverse range of cases in our curated dataset (left, right). Distribution of surgical frames by procedure type (center).
Autonomous surgery has long been a promising area of research, with the potential to revolutionize the way surgical procedures are performed [25]. To achieve this goal, significant advancements in perception and computer vision are necessary, allowing robotic systems to accurately interpret and respond to the complexities of surgical environments [38]. However, training models for surgical perception is far from trivial due to the limited availability and high quality requirements of annotated data [59, 39, 45]. The complexities of medical data and stringent regulations governing patient privacy create significant hurdles for collecting and annotating relevant information [19, 2]. The cumulative effect of all these challenges has resulted in significantly limited computer vision datasets for surgical applications, with many containing fewer than 10K images [50]. These datasets, which may suffice for validation on specific computer vision tasks and surgical procedures, fall short in providing an adequate foundation for training models that can generalize effectively across diverse surgical settings and datasets.
Recent advances in self-supervised learning have opened up new research avenues, pretrained foundation models can significantly expedite progress, potentially reducing the need for annotated data by utilizing large-scale unlabeled datasets [49]. Despite this progress, ongoing data collection challenges persist, motivating the exploration of two key aspects: the sufficiency of publicly available surgical videos as a foundation for building a comprehensive dataset, and the potential of self-supervised models trained on this dataset to achieve state-of-the-art performance in surgical downstream tasks. By addressing these aspects, we aim to provide valuable insights into the feasibility and efficacy of using public domain data for training machine learning models in the field of surgical video analysis.
@@
Figure 3: Data curation pipeline. (a) Surgical videos are collected from YouTube. (b) The videos are summarized as storyboards (trivial to annotate and classify). We train a storyboard classifier to identify videos with rich surgical content, and manually verify the selected videos. (c) During the video selection and trimming phase, a trained frame classifier categorizes each frame as either surgical or non-surgical, allowing us to trim non-surgical frame segments at the beginning and end of each video. We further filter videos by requiring at least 90% of their frames to be surgical frames. (d) In the video preprocessing stage, we obliterate non-surgical frames. (e) During the video annotation phase, we utilize the video title as a primary cue to determine the procedure and surgery type. In cases where the video titles do not explicitly include any of our procedure name keywords (Fig. 2), we employ ChatGPT to match the titles with surgical procedure types. The set of final videos, and their corresponding labels, are manually quality controlled.
Contributions. We propose a novel data curation pipeline (Fig. 3) for collecting videos of selected surgical procedures, capturing exclusively surgical footage while discarding non-surgical content. Building upon this data curation pipeline, we present Surg-3M, a dataset for computer vision applications in surgery. Our Surg-3M dataset is large and diverse, comprising more than 4K surgical videos totaling 938 hours of high-quality footage (Table 1). To the best of our knowledge, Surg-3M is the largest open-access surgical video dataset to date, 18×18$\times$18 × larger than the largest published surgical dataset [54]. To model the features of surgical video frames, we provide SurgFM, a pretrained model based on the ConvNeXt architecture [36] and DINO [8]. Additionally, we propose a novel augmented distillation method (Fig. 4) designed to enhance the local-to-global self-distillation mechanism described in  [8]. SurgFM outperformed the current state-of-the-art models (Fig. 1) on Cholec80 [54], M2CAI16 [54, 53], AutoLaparo [57], and CholecT50 [44]. Additionally, it helps to significantly reduce the annotation effort (Fig. 5). Building on our Surg-3M annotations, we introduce two new supervised learning applications: multi-label surgical procedure classification and surgery type classification. Leveraging our image foundation model, SurgFM, we propose SurgFM-Vid, a video classifier that achieves top performance on these new tasks, serving as a baseline for future research.
@@
Table 1: Comparison of Surg-3M’s size to other existing public surgical datasets. The number of frames is obtained by sampling videos at one frame per second.
Dataset # Procedures # Videos # Hours # Frames # Frames (1 fps)
M2CAI16 [54] 1 41 $26$ $2.4$M $95$K
Cholec80 [54] 1 80 $51$ $4.6$M $183$K
CholecT50 [44] 1 50 $28$ $2.5$M $101$K
AutoLaparo [57] 1 21 $23$ $2.1$M $83$K
SAR-RARP50 [48] 1 50 $45$ $9.8$M $163$K
GraSP [4] $1$ $13$ $33$ $3.5$M $117$K
Surg-3M (ours) $\mathbf{35}$ $\mathbf{4194}$ $\mathbf{938}$ $\mathbf{85}$M $\mathbf{3.4}$M
@@ 2 Related Work
Surgical datasets. Previous works have released open-access datasets aimed at addressing specific surgical downstream tasks and procedures (Table 1). Despite their valuable contributions, these open-access datasets suffer from a narrow scope of surgical procedures and limited dataset sizes. Besides these open-access datasets, there are several private surgical datasets, including SVL [63], with 1.4K videos, and SAGES CVS [47], with 1K videos. However, the restricted accessibility of these datasets constrains their utility and hinders their potential for more widespread use. To overcome these limitations, our work introduces Surg-3M. Compared with Cholec80 [54], the largest open-access surgical dataset, Surg-3M provides 18$\times$ more surgical images and 35$\times$ the surgical procedure types.
Self-supervised learning. Self-supervised learning (SSL) has gained significant attention due to its ability to reduce reliance on large-scale annotated datasets, a benefit that is particularly relevant for surgical applications. Recently, Ramesh et al. [49] investigated the effectiveness of different SSL approaches for surgical applications, including MoCov2 [11], SwAV [7], SimCLR [10], and DINO [8]. Their experiments showed that DINO consistently outperformed other SSL methods on both surgical phase recognition and surgical tool presence detection tasks, highlighting its potential in surgical applications.
@@
As part of a broader effort to advance self-supervised surgical models, Hirsch et al. [24] built a private surgical dataset and trained Masked Siamese Networks (MSN) [3] on this dataset, achieving near-state-of-the-art results. While the research presented in this study shows promising results, its broader applicability and reproducibility are somewhat constrained by the lack of public availability of the underlying dataset, as well as a relatively narrow evaluation scope that only considers performance on two specific task-dataset pairs: surgical phase recognition on Cholec80 [54], and polyp characterization on PolypSet [32]. Similarly, in Endo-FM [58], the authors proposed to train a student-teacher foundation model on a composite dataset that combined private and public data sources. Although showing encouraging results, the approach is also of limited accessibility. More recently, Batic et al. [5] created a dataset of 70K images merging nine public datasets. They trained a model based on masked autoencoders (MAE)  [23], but showed limited performance, likely due to their reduced dataset size.
[…]
Method Phase recognition (Cholec80) Tool detection (Cholec80) Action recognition (CholecT50)
Accuracy (%) F1-score (%) mAP (%) F1-score (%) mAP (%) F1-score (%)
DINO (frozen) [8] $69.8\pm 1.5$ $62.5\pm 1.5$ $67.8\pm 3.1$ $59.3\pm 4.0$ $36.1\pm 1.4$ $28.9\pm 2.7$
DINOv2 (frozen) [46] $67.0\pm 1.7$ $60.3\pm 1.8$ $56.1\pm 3.4$ $41.2\pm 2.6$ $27.2\pm 2.2$ $17.2\pm 0.8$
MAE (frozen) [23] $55.2\pm 1.4$ $49.6\pm 1.2$ $40.4\pm 1.4$ $22.9\pm 0.9$ 20.7 $\pm$ 1.3 13.4 ±$\pm 1.3$
[…]
3 Proposed Surg-3M Dataset
@@
To develop a robust surgical foundation model, it is essential to create a comprehensive, large-scale dataset of diverse surgical images that covers the whole spectrum of minimally invasive surgical procedures executed within the operating room. To fulfill this need we propose Surg-3M, a dataset comprising more than 4K videos totaling 938 hours of surgical footage covering 35 distinct procedures, including videos from robotic-assisted surgeries and non-robotic traditional laparoscopies (more details on these surgical techniques are provided in Sec. A of the supplementary material).
Ground truth and proposed new tasks. Each of the 4194 surgical videos is annotated for two tasks: 1) multi-label procedure type classification, where the objective is to identify the procedure[s] performed in an intervention’s video from among 35 possible options (Fig. 2); and 2) binary surgery type classification, where the goal is to indicate whether a surgery is conducted robotically or non-robotically. The rationale for these annotations is two-fold. Firstly, it allows us to validate the comprehensiveness of our dataset by ensuring that it encompasses a broad spectrum of surgical procedures. Secondly, it facilitates the development of new methods that use this information as a key input (e.g., our SurgFM foundation model, video classification/retrieval, video similarity measurement, procedure-specific estimation of surgical remaining time).
Scenes in surgical videos unfold at a relatively slow pace, as surgeons must ensure patient safety. Consequently, we extract frames from Surg-3M at 1 fps for our foundation model pretraining, following the standard practice in the field of surgical computer vision [54, 57, 44, 48, 53]. In the following sections we detail how we collected, curated, and annotated these videos. An overview of the curation pipeline is shown in Fig. 3.
@@ 3.1 Dataset Curation
Video collection. Inspired by works in other fields [15, 9, 61, 1, 40], we collected our videos from YouTube. For each of the $35$ procedures in Fig. 2, we searched for videos of their robotic and non-robotic variants using the strings “robotic $<$procedure type$>$” and “laparoscopic $<$procedure type$>$”, respectively. We run the video collection process until we had $500$ different videos collected for each procedure leading to a total of $\approx$18K raw videos.
@@
Video classification. The raw videos collected based on search keywords contained a diverse range of content, including patient testimonials and conference presentations. To avoid including non-surgical videos in our dataset, we classified the videos as surgical and non-surgical (Fig.  3b). This pipeline consisted of three stages. First, a dataset of 2160 surgical and 1910 non-surgical video storyboards (Fig. 3b) was annotated. Second, a binary video storyboard classifier was trained on this dataset and run over all the raw videos. Finally, manual verification was performed on the videos classified as surgical.
@@
Video selection and trimming. A video classified as surgical may still contain non-surgical frames (e.g., presentation slides). Therefore, we trained a surgical frame classifier and used it to 1) trim the non-surgical content present at the beginning and end of videos, and 2) discard videos containing more than 10% of the frames estimated as non-surgical after trimming (Fig. 3c).
Video preprocessing. We use our previously trained frame classifier to detect and obliterate those frames that are estimated to be non-surgical within the trimmed videos. After this process, we manually quality control the videos to verify that their frames are surgical.

### v3

# LEMON: A Large Endoscopic MONocular Dataset and Foundation Model for Perception in Surgical Settings

@@ Abstract
Traditional open-access datasets focusing on surgical procedures are often limited by their small size, typically consisting of fewer than 100 videos and less than 30 hours of footage, which leads to poor model generalization. To address this constraint, a new dataset called LEMON has been compiled using a novel aggregation pipeline that collects high-resolution videos from online sources. Featuring an extensive collection of over 4K surgical videos totaling 938 hours (85 million frames) of high-quality footage across multiple procedure types, LEMON offers a comprehensive resource surpassing existing alternatives in size and scope, including two novel downstream tasks. To demonstrate the effectiveness of this diverse dataset, we introduce LemonFM, a foundation model pretrained on LEMON using a novel self-supervised augmented knowledge distillation approach. LemonFM consistently outperforms existing surgical foundation models across four downstream tasks and six datasets, achieving significant gains in surgical phase recognition (+9.5pp, +9.4pp, and +8.4pp of Jaccard in AutoLaparo, M2CAI16, and Cholec80), surgical action recognition (+4.4pp of mAP in CholecT50), surgical tool presence detection (+5.3pp and +10.2pp of mAP in Cholec80 and GraSP), and surgical semantic segmentation (+8.3pp of mDice in CholecSeg8k). LEMON and LemonFM will serve as foundational resources for the research community and industry, accelerating progress in developing autonomous robotic surgery systems and ultimately contributing to safer and more accessible surgical care worldwide. Our dataset, models, and code are publicly available at https://github.com/visurg-ai/LEMON.
@@ 1 Introduction
Figure 1: Performance comparison between LemonFM, our foundation model pretrained on LEMON, and other surgical foundation models. The numeric labels mark the axis ticks. The full results are available in Table 3 and 4.
@@
Figure 2: Data curation pipeline. (a) Surgical videos are collected from YouTube. (b) The videos are summarized as storyboards (trivial to annotate and classify). We train a storyboard classifier to identify videos with rich surgical content, and manually verify the selected videos. (c) During the video selection and trimming phase, a trained frame classifier categorizes each frame as either surgical or non-surgical, allowing us to trim non-surgical frame segments at the beginning and end of each video. We further filter videos by requiring at least 90% of their frames to be surgical frames. (d) In the video preprocessing stage, we obliterate non-surgical frames and non-surgical regions within surgical frames. (e) During the video annotation phase, we utilize the video title as a primary cue to determine the procedure and surgery type. In cases where the video titles do not explicitly include any of our procedure name keywords, we employ ChatGPT to match the titles with surgical procedure types. Both the final videos and their corresponding labels are manually quality controlled.
Autonomous surgery has long been a promising area of research, with the potential to revolutionize the way surgical procedures are performed [24]. It promises safer and more consistent operations, fewer complications, shorter operative times, lighter workloads for surgeons, and wider access to high‑quality care. To achieve this goal, significant advancements in perception and computer vision are necessary, allowing robotic systems to accurately interpret and respond to the complexities of surgical environments [38]. However, training models for surgical perception is far from trivial due to the limited availability and high quality requirements of annotated data [54, 40, 44]. The complexities of medical data and stringent regulations governing patient privacy create significant hurdles for collecting and annotating relevant information [17, 2]. The cumulative effect of all these challenges has resulted in significantly limited computer vision datasets for surgical applications, with many containing fewer than 10K images [47]. These datasets, which may suffice for validation on specific computer vision tasks and surgical procedures, fall short in providing an adequate foundation for training models that can generalize effectively across diverse surgical settings and datasets.
Recent advances in self-supervised learning have opened up new research avenues. Pretrained foundation models can significantly expedite progress, potentially reducing the need for annotated data by utilizing large-scale unlabeled datasets [46]. Despite this progress, ongoing data collection challenges persist, motivating the exploration of two key aspects: the sufficiency of publicly available surgical videos as a foundation for building a comprehensive dataset, and the potential of self-supervised models trained on this dataset to achieve state-of-the-art performance in surgical downstream tasks. By addressing these aspects, we aim to provide valuable insights into the feasibility and efficacy of using public domain data for training machine learning models in the field of surgical vision analysis.
@@
Table 1: Comparison of LEMON’s size to other existing public surgical datasets.
Dataset Procedure Video Hour Frame Frame (1 fps)
M2CAI16 1 41 262626 2.42.42.4M  [^95] K
Cholec80 1 80 515151 4.64.64.6M 183183183K
CholecT50 1 50 282828 2.52.52.5M 101101101K
AutoLaparo 1 21 232323 2.12.12.1M  [^83] K
GraSP 111 131313 333333 3.53.53.5M 117117117K
GenSurgery 282828 318231823182 680680680  [^70] M 2.42.42.4M
SurgeNetXL 232323 325332533253 576576576  [^52] M 2.12.12.1M
LEMON $\mathbf{35}$ $\mathbf{4194}$ $\mathbf{938}$ $\mathbf{85}$M $\mathbf{3.4}$M
Figure 3: Comparison between LEMON and other surgical datasets sourced from the web. Video segments from each dataset illustrating the data curation differences (left). Procedures which are not covered by both GenSurgery and SurgeNetXL or any other public dataset (right).
Contributions. To address the scarcity of high‑quality surgical data and the practical difficulty of collecting it, we propose a novel multi-stage data curation pipeline designed specifically for online surgical videos (Fig. 2). In contrast to previous efforts that relied on limited, manually curated validation processes with datasets containing non-surgical content, our pipeline employs a systematic approach to filter out non-surgical videos, remove non-surgical content from surgical videos, and verify the curated surgical videos and their annotations. Building upon this data curation pipeline, we present LEMON, a large-scale dataset comprising more than 4K surgical videos totaling 938 hours (85 million frames) of high-quality footage (Table 1). Additionally, we propose two novel supervised video classification tasks based on the accompanying annotations: multi-class surgical procedure classification and binary surgery type classification, and establish a leaderboard for them. To the best of our knowledge, LEMON is the largest open access surgical dataset to date, both in total frames and videos and in the breadth of distinct surgical procedures it includes. To model the features of surgical frames, we introduce LemonFM, a foundation model pretrained on LEMON that leverages self-supervised learning with our proposed augmented knowledge distillation method. We evaluated LemonFM on six datasets: AutoLaparo [52], Cholec80 [51], M2CAI16 [51], GraSP [4], CholecT50 [43], and CholecSeg8k [23], covering four tasks including phase recognition, tool presence detection, action recognition, and semantic segmentation. Under both linear probing and full fine-tuning protocols, LemonFM consistently surpasses existing surgical foundation models and task-specific architectures (Fig. 1, Table 2, 3, 4). Moreover, LemonFM retains these advantages even when fine‑tuned with only half of the labeled data. These results showcase LemonFM’s strong generalization and data efficiency over diverse surgical scenarios and underscore the value of LEMON as a broadly useful resource for the community. We provide open access to our dataset, models, and code for data curation and training.
@@ 2 Related Work
Surgical datasets. Previous works have released open-access datasets aimed at addressing specific surgical vision tasks and procedures (Table 1). Despite valuable, these datasets suffer from a narrow scope of surgical procedures and limited dataset sizes. GenSurgery [48] and SurgeNetXL [25] expand procedural coverage by compiling public videos, but the lack of a curation step to remove the non-surgical content from the public videos (Fig. 3) might introduce spurious features that can confound discriminative models. Additionally, several private datasets exist, such as SVL [57] and SAGES CVS [45], but their restricted access limits their utility and broader adoption. To overcome these limitations, we introduce a novel data curation pipeline that yields LEMON, the largest open-access surgical dataset to date, surpassing previous datasets in both the number of frames and videos and the diversity of distinct surgical procedures represented.
@@
Self-supervised learning. Self-supervised learning (SSL) has gained significant attention due to its ability to reduce reliance on large-scale annotated datasets, a benefit that is particularly relevant for surgical applications. Recently, Ramesh et al. [46] investigated the effectiveness of different SSL approaches for surgical applications. Their experimental results show that DINO [8] outperformed MoCov2 [11], SwAV [7], and SimCLR [10] in surgical applications. As part of a broader effort to advance self-supervised surgical models, Hirsch et al. [22] built a private surgical dataset and trained a Masked Siamese Networks (MSN) [3] on it. Despite encouraging results, the work’s wider applicability and reproducibility are limited by its reliance on private data and a relatively narrow evaluation scope: surgical phase recognition on Cholec80 and polyp characterization on PolypSet [33]. Similarly, in Endo-FM [53], the authors proposed to train a student-teacher foundation model on a composite dataset that combined private and public data sources. Although showing promising results, the approach is also of limited accessibility. More recently, Batic et al. [5] created a dataset by merging nine public datasets. They trained a model based on masked autoencoders (MAE) [21], but showed limited performance, likely due to their reduced dataset size.
3 Proposed LEMON Dataset
@@
To develop a robust surgical foundation model, it is essential to create a comprehensive, large-scale dataset of diverse surgical images that covers the whole spectrum of minimally invasive surgical procedures executed within the operating room. To fulfill this need, we propose LEMON, a dataset comprising more than 4K videos totaling 938 hours of surgical footage that covers 35 distinct procedures, including videos from robotic-assisted surgeries and non-robotic traditional laparoscopies (see more details on these surgical techniques in the supplementary material).
@@ 3.1 Dataset Curation
We detail how we collected, curated, and annotated these videos. An overview of the pipeline is shown in Fig. 2.
@@
Video collection. We first compiled a list of 35 distinct minimally invasive surgical procedure types [39] (see Fig. 1 in supplementary material). Inspired by works in other fields [14, 9, 56, 1, 41], we collected our videos from YouTube. For each of the 353535 procedures, we searched for videos of their robotic and non-robotic variants using the strings “robotic $<$procedure type$>$” and “laparoscopic $<$procedure type$>$”, respectively. We run the video collection process until we had 500500500 different videos collected for each procedure leading to a total of $\approx$18K raw videos.
@@
Video classification. The raw videos collected based on search keywords contained a diverse range of content, including patient testimonials and conference presentations. To avoid including non-surgical videos in our dataset, we classified the videos as surgical and non-surgical (Fig.  2b). First, we created 4$\times$4-image video storyboards (i.e., a single image containing 16 key video frames) for each video using the method described in [16]. Second, a dataset of 2160 surgical and 1910 non-surgical video storyboards (Fig. 2b) was annotated. Third, a binary video storyboard classifier (i.e., a ResNet18 [20]) was trained on this dataset and run over all the raw videos. Finally, manual verification was performed on the videos classified as surgical.
@@
Video selection and trimming. A video classified as surgical may still contain non-surgical frames (e.g., introductory and conclusion slides). Experimentally, we found that the start and end of the surgical footage can be reliably identified by finding the first and last three consecutive surgical frames (sampling the video at 1 fps). Thus, we manually annotated a dataset of 7967 frames (5481 surgical and 2486 non-surgical), trained a surgical frame classifier (i.e., a ResNet18) on it, and used this model to: 1) trim the non-surgical content present at the beginning and end of videos, and 2) discard videos containing more than 10% of the frames estimated as non-surgical after trimming (Fig. 2c).
Video preprocessing. As the videos were from YouTube, we cannot assume that all frames within the trimmed videos are entirely free from non-surgical content (e.g., intraoperative out-of-body views, medical annotations, user-interface details, equipment logos). Therefore, we used our trained frame classifier to detect and obliterate the intraoperative non-surgical frames (Fig. 2d). Additionally, we manually annotated 2719 surgical frames with 4584 non-surgical bounding-box instances and trained a YOLOv8 model [31] to detect and obliterate the non-surgical content in surgical frames. After this process, we manually quality controlled all the remaining curated videos.